### v1

Abstract
Phase retrieval from phaseless short-time Fourier transform (STFT) measurements is known to be inherently unstable when measurements are taken with respect to a single window. While an explicit inversion formula exists, it is useless in practice due to its instability. In this paper, we overcome this lack of stability by presenting two multi-window approaches that rely on a “good coverage” of the time-frequency plane by the ambiguity functions of the windows. The first is to use the fractional Fourier transform of a dilated Gauss function with various angles as window functions. The essential support of a superposition of the ambiguity function from such window functions is of a “daffodil shape”, which converges to a large disc as more angles are used, yielding a much broader coverage in the time-frequency domain. The second approach uses Hermite functions of various degrees as the window functions. The larger the degree, the wider the ambiguity function but with zeros on circles in the time-frequency domain. Combining Hermite functions of different degrees, we can achieve a wide coverage with zeros compensated by the essential support of the ambiguity function from other Hermite windows. Taking advantage of these multi-window procedures, we can stably perform STFT phase retrieval using the direct inversion formula.
keywords:
phase retrieval, short-time Fourier transform, stability, direct method
{MSCcodes}
42-08, 42C99, 45Q05, 49K40
1 Introduction
In many scientific and engineering fields, seeking to reconstruct a signal or image from magnitude measurements alone is a common challenge. In these applications, only the absolute values of a set of linear measurements can be acquired so that phase information is missing. In some instances, the phase cannot be directly measured or the phase measurements are too noisy to be of use. Phase retrieval addresses this problem by using computational algorithms that estimate the missing phase data, enabling the full reconstruction of the original signal or object. This technique is essential in areas such as optics, crystallography, and diffraction imaging, where direct phase measurement is often impractical. Specific audio processing applications also require phase retrieval from certain magnitude-only measurements.
In this paper, we focus on a specific instance of phase retrieval, which aims at recovery from magnitude measurements of the short-time Fourier transform (STFT), that we will first introduce: For a given window function $g\in L^{2}(\mathbb{R})$, the linear mapping that takes any function $f\in L^{2}(\mathbb{R})$ to its short-time Fourier transform (STFT) or windowed Fourier transform is defined as
(1) $V_{g}f(x,y):=\int_{\mathbb{R}}f(t)\overline{g(t-x)}e^{-2\pi ity}\,\mathrm{d}t.$
The special case of the STFT with a Gauss window $g(t)=\varphi(t):=e^{-\pi t^{2}}$ is sometimes referred to as the Gabor transform. The STFT enjoys many nice properties such as the fact that it is a linear isometry from $L^{2}(\mathbb{R})$ to $L^{2}(\mathbb{R}^{2})$ up to the factor $\|g\|_{L^{2}(\mathbb{R})},$ i.e.,
$\|V_{g}f\|_{L^{2}(\mathbb{R}^{2})}=\|f\|_{L^{2}(\mathbb{R})}\|g\|_{L^{2}(\mathbb{R})}.$
Thus, the reconstruction from the STFT is unique and stable, since for any $f_{1},f_{2}\in L^{2}(\mathbb{R})$:
$\|V_{g}f_{1}-V_{g}f_{2}\|_{L^{2}(\mathbb{R}^{2})}=\|V_{g}(f_{1}-f_{2})\|_{L^{2}(\mathbb{R}^{2})}=\|f_{1}-f_{2}\|_{L^{2}(\mathbb{R})}\|g\|_{L^{2}(\mathbb{R})}.$
Furthermore, an explicit inversion formula for the STFT can be given [18].
A question well motivated by applications in audio processing and ptychographic imaging is that of recovering $f$ from partial information of its STFT. More precisely, in these applications, one is interested in recovering a signal $f$ from the magnitude of its STFT, i.e., from $\left|V_{g}f\right|$. This is a special case of a phase retrieval problem, also referred to as STFT (or Gabor) phase retrieval. Note that this is a nonlinear recovery problem. While it is not straightforward to resolve the missing phase information, there is, interestingly, a formula that relates $\left|V_{g}f\right|$ to the ambiguity function $\mathcal{A}f$ of $f$, defined as
$\mathcal{A}f(x,y)=\int_{\mathbb{R}}f\left(t+x/2\right)\overline{f\left(t-x/2\right)}e^{-2\pi ity}\,\mathrm{d}t,$
which can be more compactly written as
$\mathcal{A}f(x,y)=e^{\pi ixy}\,V_{f}f(x,y).$
Its relation with the magnitude of the STFT of $f$ is given by (see e.g. [24])
(2) $\mathcal{F}\left(\left|V_{g}f\right|^{2}\right)(y,-x)=\mathcal{A}f(x,y)\cdot\overline{\mathcal{A}g(x,y)}=V_{f}f(x,y)\cdot\overline{V_{g}g(x,y)}\,,$
where $\mathcal{F}$ denotes the two-dimensional Fourier transform on $L^{2}(\mathbb{R}^{2})$, and $\cdot$” denotes pointwise multiplication. We note that the above equation is a direct consequence of the orthogonality relations of the STFT (cf. [16, Theorem 3.2.1.]). Thus, at least in theory, the ambiguity function $\mathcal{A}f$ can be recovered from $\left|V_{g}f\right|$, whenever $\mathcal{A}g$ is non-zero, through
(3) $\mathcal{A}f(x,y)=\frac{\mathcal{F}\left(\left|V_{g}f\right|^{2}\right)(y,-x)}{\overline{\mathcal{A}g(x,y)}}.$
We note that if a value $c\in\mathbb{R}$ is known, for which $f(c)\neq 0,$ then $f$ can be reconstructed from its ambiguity function via a one-dimensional inverse Fourier transform with respect to the second variable and setting $t=x/2+c:$
(4) $f(x+c)=\frac{1}{\overline{f(c)}}\int_{\mathbb{R}}\mathcal{A}f(x,y)\,e^{2\pi i(\frac{x}{2}+c)y}\mathrm{d}y,\quad x\in\mathbb{R},$
or equivalently,
$f(x+c)=\frac{1}{\overline{f(c)}}\int_{\mathbb{R}}V_{f}f(x,y)\,e^{2\pi i(x+c)y}\mathrm{d}y,\quad x\in\mathbb{R}.$
In practice, however, this approach is very unstable in general: if the window $g$ is well-localized in either time or frequency, which is the case for meaningful windows, this will be reflected in its ambiguity function $\mathcal{A}g$. More precisely, if, for example, $g$ decays fast in time, then $\mathcal{A}g(x,y)$ will also decay rapidly in the $x-$direction. The same is true for fast decay of $g$ in frequency, yielding a corresponding decay of $\mathcal{A}g$ in $y$.
Note that for many applications, it is desirable to work with a window that is well localized in time and frequency because, in this case, the STFT yields a representation of the signal that is better localized in time and frequency, i.e., it can better describe which frequencies occurred (roughly) at what time (with the limit being the uncertainty principle). Hence, even if $\mathcal{A}g$ is nonzero everywhere on $\mathbb{R}^{2}$, its magnitude typically decays very fast, making recovery of $\mathcal{A}f$ via Equation (3) highly unstable. Take for example the Gauss window $\varphi$. Its ambiguity function is a two-dimensional Gauss function, i.e.,
$\mathcal{A}\varphi(x,y)=\frac{1}{\sqrt{2}}e^{-\frac{\pi}{2}(x^{2}+y^{2})},$
which decays exponentially as $x,y\rightarrow\pm\infty$. Recovery of $\mathcal{A}f$ via pointwise inversion by $\mathcal{A}\varphi$, as Equation (3) suggests, is therefore numerically infeasible, unless the signal $f$ itself also happens to be very well-localized in time and in frequency.
The goal of this paper is to propose algorithms that can employ the convenient and explicit inversion formula (3) in a way that is numerically stable. Existing reconstruction algorithms are much more involved: the most prominent algorithms employed for a wide range of phase retrieval problems are iterative projection algorithms, such as the Gerchberg–Saxton algorithm [15] and Fienup’s algorithm [14], Averaged Alternating Reflections [7] and its variants [11, 13, 21, 20, 10]. A more recent popular iterative scheme is Wirtinger Flow [8]. There also exist methods based on convex relaxation, such as PhaseLift [9], which is based on the idea of lifting the problem to higher dimensions and then recovering a low-rank solution of a linear problem instead of solving the non-linear phase retrieval problem. In practice, however, these algorithms have not been adopted as the iterative methods mentioned before are computationally much more efficient.
Another method, specific to STFT phase retrieval, is phase gradient heap integration (PGHI), which recovers the phase of $V_{g}f$ point by point along a grid on the time-frequency plane [22]. By far, we have not described the entire literature on phase retrieval algorithms here. For a thorough description, we refer to [12].
In this paper, we propose a different approach to STFT phase retrieval. The idea is to use the simple and explicit inversion formula (3) but with measurements with respect to several windows instead of one. This way, we can give theoretical guarantees for stability (cf. Section 3).
It is important to emphasize that not only does the specific reconstruction formula (3) exhibit poor stability properties when used with a single window function, but Gabor phase retrieval itself (from a single window function) is inherently unstable [4]. The stability properties can be linked to the connectedness of the Gabor transform magnitude as a function in the time-frequency plane [18]. Stability can be restored when global recovery of the phase is relaxed and one only aims at recovering the phase on each connected component individually [3]. In this work, we intend to take a different route: instead of drawing measurements from one single window $g,$ we take the STFT magnitudes from several different windows. This way, we overcome the trade-off between stability and disconnectedness of the measurements. We aim at stably recovering the signal globally, instead of on each connected component individually (the latter being inherent to the single window setup). The central question we address in this work is the following:
Question. Suppose we can take the STFT magnitude with respect to several windows $\{g_{j}\}_{j=1}^{N}$, obtaining a family of measurements $\{\left|V_{g_{j}}f\right|\}_{j=1}^{N}$. Can we stably recover $f$ based on the reconstruction formula (3) with a collection of properly chosen windows $\{g_{j}\}_{j=1}^{N}$?
The rest of the paper is organized to address this key question. In Section 2, we propose two multi-window algorithms for STFT phase retrieval — one utilizing fractionally Fourier transformed dilated Gaussian windows, and the other employing Hermite functions of varying degrees. Section 3 establishes mathematical results on the improved inverse problem stability provided by these two multi-window approaches. In Section 4, we present several numerical examples, and conclusions are drawn in Section 5.
2 Multi-window approaches for STFT phase retrieval
In what follows, we propose a framework to use Equation (3) constructively, via employing multiple windows $\{g_{j}\}_{j=1}^{N}$ with different time-frequency concentration each. This way, each set of phaseless measurements $\left|V_{g_{j}}f\right|$ for a given $j$ allows for stable extraction of a portion of $\mathcal{A}f,$ with the region of extraction dictated by the decay of $\mathcal{A}g_{j}.$ Hence, by acquiring more measurements, we can patch together the different pieces of $\mathcal{A}f$ and globally recover $f$. Note that the ambiguity function has its maximum at the origin $(0,0)$ [16]. Therefore, we cannot create windows with ambiguity functions centered at a point different from the origin.
We propose recipes for stable STFT phase retrieval from multiple windows through the following approach:
For some fixed $\varepsilon>0$ and a family of windows $\{g_{j}\}_{j=1}^{N},$ such that $\{\left|V_{g_{j}}f\right|\}_{j=1}^{N}$ is measured, proceed as follows:
  • For each window $g_{j},$ determine the region $\Omega_{j}\subset\mathbb{R}^{2}$ such that $\left|\mathcal{A}g_{j}\right|>\varepsilon$ on $\Omega_{j}$. Choose the $\Omega_{j}$’s to be pairwise disjoint.
  • For each $j,$ recover $\mathcal{A}f$ on $\Omega_{j}$ by employing the measurements $\left|V_{g_{j}}f\right|$ in (2).
  • Combine these partial reconstructions to a global approximation of $\mathcal{A}f$.
  • Reconstruct an approximation of $f$ (up to a constant global phase factor) from the approximation of its ambiguity function $\mathcal{A}f$ through Equation (4).
In particular, we propose two architectures for $\left\{g_{j}\right\}_{j=1}^{N}$. The first is based on fractional Fourier transforms of dilated Gaussians, while the second is based on Hermite functions of different order. These choices are based on properties of the ambiguity functions of these windows. Before we describe the two measurement systems, we first record some preliminary facts that will be employed.
2.1 Dilated Gauss windows under the fractional Fourier transform
To begin with, we give the definition of the fractional Fourier transform (FrFT) $\mathcal{F}_{\alpha}$ of $f\in L^{1}(\mathbb{R})$ by the angle $\alpha$, which will be essential in what follows:
$\mathcal{F}_{\alpha}f(y):=c_{\alpha}e^{\pi iy^{2}\cot\alpha}\int_{\mathbb{R}}f(t)e^{\pi it^{2}\cot\alpha}e^{-2\pi ity/\sin\alpha}\mathrm{d}t,$
when $\𝜋\alpha=\mathbb{R}\backslash\pi\mathbb{Z}$. Here, $c_{\alpha}$ is the square root of $1-i\cot\alpha$ with positive real part. For angles $\alpha=k\pi,k\in\mathbb{Z}$, the fractional Fourier transform is defined as $\mathcal{F}_{k\pi}f:=f$ for $k$ even, and as $\mathcal{F}_{k\pi}f(y):=f(-y)$, for $k$ odd. Note that $\mathcal{F}_{\pi/2}=\mathcal{F},$ i.e., it is the standard Fourier transform. By a classical density argument, the fractional Fourier transform can be extended to all functions in $L^{2}(\mathbb{R})$ [19].
A property crucial to our first approach is the effect of the fractional Fourier transform on the STFT, which can be described as follows [6]:
(5) $V_{\mathcal{F}_{\alpha}g}\mathcal{F}_{\alpha}f(x,y)=V_{g}f(R_{\alpha}(x,y))e^{\pi i\sin\alpha\left((x^{2}-y^{2})\cos\alpha-2xy\sin\alpha\right)},\quad x,y\in\mathbb{R},$
for $\alpha\in\mathbb{R}$ and $f,g\in L^{2}(\mathbb{R}),$ where $R_{\alpha}(x,y)=(x\cos\alpha-y\sin\alpha,x\sin\alpha+y\cos\alpha)$, i.e., it is the rotation of the vector $(x,y)\in\mathbb{R}^{2}$ by the angle $\alpha.$ Thus,
$\left|V_{\mathcal{F}_{\alpha}g}\mathcal{F}_{\alpha}f(x,y)\right|=\left|V_{g}f(R_{\alpha}(x,y))\right|,\quad\mbox{ for }x,y\in\mathbb{R},$
so that $\left|V_{\mathcal{F}_{\alpha}g}\mathcal{F}_{\alpha}f\right|$ is a rotation by $\alpha$ of $\left|V_{g}f\right|$.
Let us now consider Gauss functions parameterized by a dilation factor $a>0,$
$\varphi^{a}(t):=a^{1/4}e^{-a\pi t^{2}},$
for which $\left\|\varphi^{a}\right\|_{L^{2}(\mathbb{R})}=2^{-1/4}.$ Computing the ambiguity function of $\varphi^{a}$ yields
$\mathcal{A}\varphi^{a}(x,y)=\frac{1}{\sqrt{2}}e^{\pi ixy}e^{-\frac{\pi\left(a^{2}x^{2}+2iaxy+y^{2}\right)}{2a}}=\frac{1}{\sqrt{2}}e^{-\frac{\pi}{2}\left(ax^{2}+\frac{1}{a}y^{2}\right)}.$
Hence, the radial decay for the standard Gauss $\varphi=\varphi^{1}$ is replaced by an elliptical decay with the $x$-axis stretched by the factor $a$. Next, let $\left\{\alpha_{j}=\pi j/N\right\}_{j=1}^{N}$ be a set of angles for which we consider the family of window functions
$\phi_{j}:=\mathcal{F}_{\alpha_{j}}\varphi^{a},\quad 1\leq j\leq N,$
for some fixed dilation factor $a>1$ and some discretization parameter $N$ such that $\{\alpha_{j}\}_{j=1}^{N}$ is an equidistant discretization of the interval $[0,2\pi]$. In view of Equation (5), this choice of windows results in ambiguity functions $\mathcal{A}\phi_{j}$ that have an elliptical exponential decay stretched by a factor $a$ along the axis $(\cos\alpha_{j},\sin\alpha_{j}).$ The superposition of these “essential supports” is shaped in the form of a daffodil and approaches a disc shape by taking $N$ larger, which is also illustrated in Figure 1.
(a) $a=1$, single window
(b) $a=2$, single window
(c) $a=2$, $8$ windows
(d) $a=10$, single window
(e) $a=10$, $4$ windows
(f) $a=10$, $40$ windows
Figure 1: The absolute value of the sum of multiple fractional Fourier transformed (dilated) Gauss windows plotted in the log scale. (a) $\mathcal{A}\varphi^{1}$, where $\varphi^{1}$ is the standard Gauss; (b) $\mathcal{A}\varphi^{2}$ for the dilated Gauss with $a=2$; (c) $\sum_{j=1}^{N}\mathcal{A}(\mathcal{F}_{\tilde{\alpha}_{j}}\varphi^{2})$ with $N=8$; (d) $\mathcal{A}\varphi^{10}$ for the dilated Gauss with $a=10$; (e) $\sum_{j=1}^{N}\mathcal{A}(\mathcal{F}_{{\alpha}_{j}}\varphi^{10})$ with $N=4$; (f) $\sum_{j=1}^{N}\mathcal{A}(\mathcal{F}_{{\alpha}_{j}}\varphi^{10})$ with $N=40$.
Note that the “leaves” of the daffodil grow in length but get narrower in width as $a$ becomes larger. Thus, to approximately cover a disc of a larger diameter, the number of windows $N$ has to grow too. We also remark that the dilated Gauss function is not the only possible choice. The main property that is exploited here, is the fact that the ambiguity functions of these (dilated) Gauss functions have no zeros and are strictly lower bounded on a predefined region (which we choose to be elliptical). A result in [17] states that the Gauss function is not the only window with zero-free ambiguity function. Another example is the one-sided exponential.
The daffodil-shaped geometry shown in Figure 1e is a very simple construction and the reconstruction can be further simplified as follows. A simple calculation yields that the ambiguity functions of $\phi_{j}$ are real-valued and positive. Here, we use the notation $R_{\alpha_{j}}(x,y)=(x(\alpha_{j}),y(\alpha_{j}))$:
$\mathcal{A}\phi_{j}(x,y)$ $=e^{\pi ixy}V_{\phi_{j}}\phi_{j}(x,y)$
$=e^{\pi ixy}e^{\pi i\sin\alpha_{j}\left((x^{2}-y^{2})\cos\alpha_{j}-2xy\sin\alpha_{j}\right)}V_{\varphi^{a}}\varphi^{a}\left(R_{\alpha_{j}}(x,y)\right)$
$=e^{\pi ix(\alpha_{j})y(\alpha_{j})}V_{\varphi^{a}}\varphi^{a}(x(\alpha_{j}),y(\alpha_{j}))$
$=\mathcal{A}\varphi^{a}(R_{\alpha_{j}}(x,y)).$
Since $\mathcal{A}\varphi^{a}$ is real-valued and positive, this is also true for $\mathcal{A}\phi_{j}$. Therefore, we can directly sum the reconstruction formula (2) over $j=1,\dots,N:$
(6) $\sum_{j=1}^{N}\mathcal{F}\left(\left|V_{\phi_{j}}f\right|^{2}\right)(y,-x)=\mathcal{A}f(x,y)\cdot\sum_{j=1}^{N}\mathcal{A}\phi_{j}(x,y)\,,$
and determine $\mathcal{A}f$ on the domain
$\Omega=\left\{(x,y)\in\mathbb{R}^{2}:\sum_{j=1}^{N}\mathcal{A}\phi_{j}(x,y)>\varepsilon\right\}$
through pointwise division in (6). We propose Algorithm 1 for the multi-window inversion using fractionally Fourier transformed Gaussian windows at various angles.
Algorithm 1 A Multiscale STFT Phase Retrieval Algorithm Based on the Fractionally Fourier Transformed Gaussian Windows.
1:Given $|V_{\mathcal{F}_{\alpha}\varphi^{a}}f|$ and $\mathcal{A}(\mathcal{F}_{\alpha}\varphi^{a})$ where $\varphi^{a}$ is the dilated Gaussian, and $\mathcal{F}_{\alpha}$ denotes the Fractional Fourier transform. We consider measurements from a collection of angles $\{\alpha_{j}\}_{j=1}^{N}$ and define $\phi_{j}=\mathcal{F}_{\alpha_{j}}\varphi^{a}$. Let $\varepsilon$ be the tolerance.
2:for $j=1$ to $N$  do
3:     Compute $\phi_{j}(x,y)=\mathcal{F}(|V_{\phi_{j}}f|^{2})$ (2D FFT) and $\widetilde{G}_{j}(x,y)=G_{j}(y,-x)$.
4:end for
5:Find $\Omega=\{(x,y):\sum_{j=1}^{N}\mathcal{A}\phi_{j}>\varepsilon\}$.
6:Define $J(x,y)=\sum_{j=1}^{N}\widetilde{G}_{j}(x,y)/\sum_{j=1}^{N}\mathcal{A}\phi_{j}(x,y)$, for $(x,y)\in\Omega$ and $J(x,y)=0$ on $\Omega^{c}$.
7:Compute $\widehat{J}(x,\cdot)=\mathcal{F}^{-1}(J(x,\cdot))$ (1D FFT).
8:Return the reconstructed signal $\widetilde{f}(x+c)=\widehat{J}(x,x/2+c)/{\overline{f(c)}}$ (assuming $f(c)\neq 0$).
A drawback of this setup is that many windows are required to achieve a good support coverage of the time-frequency plane by their ambiguity functions, such as the disc shown in Figure 1f. Next, we propose a more technical construction, but with the advantage of requiring less data in the sense that fewer windows are necessary for a good support coverage.
(a) Hermite function $h_{1}$
(b) Hermite function $h_{5}$
(c) Hermite function $h_{10}$
(d) $S_{n}^{\varepsilon}$ with $n=1$
(e) $S_{n}^{\varepsilon}$ with $n=5$
(f) $S_{n}^{\varepsilon}$ with $n=10$
(g) $\bigcup S_{n}^{\varepsilon}$ with $n=1,5$
(h) $\bigcup S_{n}^{\varepsilon}$, $n=1,5,10$
(i) $\bigcup S_{n}^{\varepsilon}$ $n=1$,$3$,$5$,$7$,$10$
Figure 2: Top row (a)-(c): the absolute value of the Hermite function ambiguity function plotted in the log scale for $n=1,5,10$ , respectively. Middle row (d)-(f): the white region reflects the stability set $S_{n}^{\varepsilon}$ defined in (8) with $\varepsilon=0.1$ for Hermite functions of degree $n=1,5,10$, respectively. Bottom row (g)-(i): the white region reflects the union of the stability set $\bigcup S_{n}^{\varepsilon}$ in (8) with $\varepsilon=0.1$ for Hermite functions with different sets of the indices $n$.
2.2 Hermite windows
For this second architecture, the set of windows will consist of Hermite (also called Hermite–Gaussian) functions of increasing order. Their ambiguity functions have rotational symmetry with a number of circular root sets related to the order of the Hermite function; see Figures 2a, 2b, and 2c. The idea of our approach is to recover $\mathcal{A}f$ as if one was “peeling an onion”: for each Hermite function, we define an annulus on which $\mathcal{A}f$ can be retrieved stably. The union of these annuli results in a large disc in the time-frequency plane. In what follows, we will make this procedure more precise. For this, we start with the definition of the Hermite function $h_{n}$ of order $n$ [1]:
Definition 2.1.
For $n\in\mathbb{N},$ the $n$-th Hermite function is defined as
$h_{n}(t):=c_{n}e^{\pi t^{2}}\frac{d^{n}}{dt^{n}}e^{-2\pi t^{2}},$
where $c_{n}$ is a constant such that $\|h_{n}\|_{L^{2}(\mathbb{R})}=1$. In particular, $h_{0}(t)=2^{1/4}e^{-\pi t^{2}}$, which is the $L^{2}-$normalized standard Gauss function.
We summarize a few basic properties of the Hermite functions in the following.
Lemma 2.2.
The Hermite functions $h_{n}$ are $L^{2}$-normalized, i.e., $\|h_{n}\|_{L^{2}(\mathbb{R})}=1$ and $h_{n}$ is the $n$-th eigenfunction of the Fractional Fourier transform with
$\mathcal{F}_{\alpha}h_{n}=e^{-i\alpha n}h_{n}.$
The family $\left\{h_{n}\right\}_{n\in\mathbb{N}}$ forms an orthonormal basis of $L^{2}(\mathbb{R})$. In particular, the Fourier transform of $h_{n}$ is $(-i)^{n}h_{n}.$
The ambiguity function of the $n$-th Hermite function is [5]
(7) $\mathcal{A}h_{n}(x,y)=e^{-\pi(x^{2}+y^{2})/2}L_{n}\left(\pi(x^{2}+y^{2})\right),$
where $L_{n}(z)$, $z\in\mathbb{R},$ is the Laguerre polynomial
$L_{n}(z):=\sum_{k=0}^{n}\frac{n!}{(n-k)!k!}\frac{(-z)^{k}}{k!}.$
We further define the $n$-th Laguerre function $\mathfrak{L}_{n}(z)$ as
$\mathfrak{L}_{n}(z)=e^{-z/2}L_{n}(z),\,\quad z\in\mathbb{R}.$
The set of Laguerre functions $\{\mathfrak{L}_{n}(z)\}_{n\in\mathbb{N}}$ forms an orthogonal basis of $L^{2}(\mathbb{R})$ and
$\mathcal{A}h_{n}(x,y)=\mathfrak{L}_{n}(\pi(x^{2}+y^{2})).$
The $n-$th Laguerre polynomial, and hence also the $n-$th Laguerre function, has exactly $n$ (strictly positive, pairwise different) roots. By Equation (7), $\mathcal{A}h_{n}$ is rotation-invariant. Hence, the root set of $\mathcal{A}h_{n}$ consists of $n$ disjoint circles, centered at the origin $(0,0)$ and thus has Lebesgue measure zero.
When solving the STFT phase retrieval problem with the help of (2), the stability of this recovery will be dictated by the absolute value of the ambiguity function: to recover the ambiguity function of the unknown signal via Equation (2) requires the division by the ambiguity function of the window. Therefore, this process is very unstable whenever the ambiguity function of the window is small in magnitude. Consequently, we need to understand when $\mathcal{A}h_{n}(x,y)$, the ambiguity function of the $n$-th Hermite window, is small in absolute value to avoid such regions in our algorithm.
To this end, we consider the stability set $S_{n}^{\varepsilon}$ for any $\varepsilon\in(0,1)$ defined by
(8) $S_{n}^{\varepsilon}:=\{(x,y)\in\mathbb{R}^{2}:|\mathcal{A}h_{n}(x,y)|>\varepsilon\}=\{(x,y)\in\mathbb{R}^{2}:|\mathfrak{L}_{n}(\pi(x^{2}+y^{2}))|>\varepsilon\}.$
In Figures 2d, 2e, and 2f, the white region reflects the set $S_{n}^{\varepsilon}$ with $\varepsilon=0.1$ and $n=1,5,10$, respectively. The stability set $S_{n}^{\varepsilon}$ is determined by the subset of $\mathbb{R}$ on which the $n$-th Laguerre function $\mathfrak{L}_{n}$ has a large absolute value. Hence, better estimates of $S_{n}^{\varepsilon}$ can aid the design of an efficient numerical strategy for phase retrieval by choosing a minimal set of Hermite windows $\{h_{n_{k}}\}$, such that $\bigcup_{k}S_{n_{k}}^{\varepsilon}$ is as large as possible. As shown in Figures 2g, 2h, and 2i, we plot the union of the stability set $\bigcup_{n\in I}S_{n}^{\varepsilon}$ for different sets of indices $I$. The larger the set $I$, the larger the union of the stability sets on which we can stably recover $\mathcal{A}f(x,y)$.
We propose Algorithm 2 for the multi-window STFT phase retrieval using Hermite functions of various degrees as the measuring windows.
Algorithm 2 A Multiscale Algorithm for the Hermite Functions
1:Given the phaseless STFT measurements $|V_{h_{n_{j}}}f|$ and the ambiguity functions $\mathcal{A}h_{n_{j}}$ where $h_{n_{j}}$ is the $n_{j}$–th Hermite function, $j=1,\ldots,N$, with $n_{1}<\cdots<n_{N}$. Let $\Omega=\emptyset$ and the tolerance be $\varepsilon>0$.
2:for $j=1$ to $N$  do
3:     Find $\Omega_{j}=\{(x,y):|\mathcal{A}h_{n_{j}}|>\varepsilon\}\setminus\Omega$ and update $\Omega\leftarrow\Omega\cup\Omega_{j}$.
4:     Compute $G_{j}(x,y)=\mathcal{F}(|V_{h_{n_{j}}}f|^{2})$ (using 2D FFT) and set $\widetilde{G}_{j}(x,y)=G_{j}(y,-x)$.
5:     Define $J(x,y)=\widetilde{G}_{j}(x,y)/\overline{\mathcal{A}h_{n_{j}}(x,y)}$, for $(x,y)\in\Omega_{j}$.
6:end for
7:Set $J=0$ on $\Omega^{c}$. Compute $\widehat{J}(x,\cdot)=\mathcal{F}^{-1}(J(x,\cdot))$ (using 1D inverse FFT).
8:Return the recovered function $\widetilde{f}(x+c)=\widehat{J}(x,x/2+c)/{\overline{f(c)}}$ (assuming $f(c)\neq 0$).
Remark 2.3.
If we are given $V_{g}g$ rather than $\mathcal{A}g$ where $g$ is the window function, we can still use Algorithms 1 and 2 with $\mathcal{A}g$ therein replaced by $V_{g}g$ and the Step 8 of both algorithms changed to $\widetilde{f}(x+c)=\widehat{J}(x,x+c)/{\overline{f(c)}}$, assuming $f(c)\neq 0$.
3 Stability estimates
The improved stability properties of our approach, evidenced by numerical examples in Section 4, can be quantified more explicitly. We first derive an estimate for Algorithm 1. A simple modification, outlined in Section 3.2, implies a similar result for Algorithm 2.
For this, suppose that on each region $\Omega_{j}\subset\mathbb{R}^{2}$ we compare a set of exact measurements
$\left\{F_{j}:=\left|V_{g_{j}}f\right|^{2}\right\}_{j=1}^{N}\,,$
and measurements corrupted by additive noise
$\left\{F_{j}^{\eta}:=\left|V_{g_{j}}f\right|^{2}+\eta_{j}\right\}_{j=1}^{N}\,,$
where $\{g_{j}\}_{j=1}^{N}$ is a set of $N$ distinct window functions. The goal is to analyze the stability of reconstructing $f$ from $\left\{F_{j}\right\}_{j=1}^{N}$ via formulas (3) and (4), where we employ (3) with $F_{j},F_{j}^{\eta}$ on $\Omega_{j}$ for each $j=1,\dots,N$.
There are two error terms that enter in the analysis: the reconstruction error due to the fact that the ambiguity function will only be determined on the bounded domain $\Omega:=\bigcup_{j}\Omega_{j}$ and the propagation of the noise terms $\eta_{j}$. We will denote by $\widetilde{f}$ and $\widetilde{f\,}^{\eta}$ the reconstructions from $F_{j}$’s and $F^{\eta}_{j}$’s, respectively. The two error terms we need to bound are on the right-hand side (RHS) of
$\left\|f-\widetilde{f\,}^{\eta}\right\|_{L^{p}(\mathbb{R})}\leq\left\|f-\widetilde{f}\right\|_{L^{p}(\mathbb{R})}+\left\|\widetilde{f}-\widetilde{f\,}^{\eta}\right\|_{L^{p}(\mathbb{R})},$
where the first term on the RHS is the approximation error and the second is the propagation of data noise.
3.1 Stability of Algorithm 1
This subsection analyzes the stability of Algorithm 1.
3.1.1 Propagation of data noise
We start by estimating the second term, $\|\widetilde{f}-\widetilde{f\,}^{\eta}\|_{L^{p}(\mathbb{R})}$. For this, let $A$ and $A^{\eta}$ denote the functions obtained on $\Omega=\bigcup_{j=1}^{N}\Omega_{j}$ through the reconstruction formula (6) from $F_{j},$ $F_{j}^{\eta}$, respectively. That is,
$A(x,y)=\frac{\sum_{j=1}^{N}\mathcal{F}\left(F_{j}\right)(y,-x)}{\sum_{j=1}^{N}\mathcal{A}\phi_{j}(x,y)},\quad A^{\eta}(x,y)=\frac{\sum_{j=1}^{N}\mathcal{F}\left(F_{j}^{\eta}\right)(y,-x)}{\sum_{j=1}^{N}\mathcal{A}\phi_{j}(x,y)}\,,\quad(x,y)\in\Omega.$
Since the reconstructions are only computed on $\Omega,$ we simply set $A$ and $A^{\eta}$ to zero on $\mathbb{R}^{2}\backslash\Omega.$ This does not affect the data noise propagation, but the fact that the ambiguity function will only be recovered on $\Omega$ will enter in the reconstruction error term later. To this end, we introduce the mixed-norm $L^{p,1}$ space consisting of all functions $U$ on $\mathbb{R}^{2}$ for which
$\left\|U\right\|_{L^{p,1}(\mathbb{R}^{2})}:=\int_{\mathbb{R}}\left(\int_{\mathbb{R}}\left|U(x,y)\right|^{p}\mathrm{d}x\right)^{1/p}\mathrm{d}y<\infty\,.$
We derive the following bounds on the propagation of data noise in the ambiguity functions.
Lemma 3.1.
Let $A,A^{\eta}$ be the reconstructions obtained by (6) from $F_{j},$ $F_{j}^{\eta},$ respectively, $j=1,\dots,N$ and for some $\varepsilon>0$, let $\Omega$ be chosen such that $\sum_{j=1}^{N}\mathcal{A}\phi_{j}(x,y)>\varepsilon$ for all $(x,y)\in\Omega$. Then,
(9) $\left\|A-A^{\eta}\right\|_{L^{p,1}(\mathbb{R}^{2})}$ $\leq\frac{C_{p}}{\varepsilon}\left\|\sum_{j=1}^{N}\eta_{j}\right\|_{W^{1,1}(\mathbb{R}^{2})},$
where $C_{p}:=\left\|\mathds{1}_{\Omega}/\omega\right\|_{L^{p,1}(\mathbb{R}^{2})}$, $\mathds{1}_{\Omega}$ denotes the characteristic function on $\Omega$ and $\omega(x,y):=1+2\pi i(x+y)$. Furthermore,
(10) $\left\|A(0,\cdot)-A^{\eta}(0,\cdot)\right\|_{L^{1}(\mathbb{R})}$ $\leq\frac{C_{\mathbb{R},\Omega}}{\varepsilon}\cdot\left\|\sum_{j=1}^{N}\eta_{j}\right\|_{L^{1}(\mathbb{R}^{2})},$
where $C_{\mathbb{R},\Omega}:=|\Omega_{x}|$ where $\Omega_{x}=\{x\in\mathbb{R}:(x,0)\in\Omega\}\subset\mathbb{R}$.
Proof 3.2.
Formula (6) gives
$\left\|A-A^{\eta}\right\|_{L^{p,1}(\mathbb{R}^{2})}=\left\|\mathds{1}_{\Omega}\frac{\sum_{j=1}^{N}T\mathcal{F}(F_{j}-F_{j}^{\eta})}{\sum_{j=1}^{N}\mathcal{A}\phi_{j}}\right\|_{L^{p,1}(\mathbb{R}^{2})}\leq\frac{C_{p}}{\varepsilon}\left\|\omega\,T\sum_{j=1}^{N}\mathcal{F}(F_{j}-F_{j}^{\eta})\right\|_{L^{\infty}(\mathbb{R}^{2})},$
where $T(\cdot)(x,y)=(\cdot)(y,-x)$. This can be further estimated by
$\left\|A-A^{\eta}\right\|_{L^{p,1}(\mathbb{R}^{2})}$ $\leq\frac{C_{p}}{\varepsilon}\left\|\widetilde{\omega}\,\mathcal{F}\left(\sum_{j=1}^{N}\eta_{j}\right)\right\|_{L^{\infty}(\mathbb{R}^{2})}$
$\leq\frac{C_{p}}{\varepsilon}\left\|\mathcal{F}\left(\sum_{j=1}^{N}(\eta_{j}-\frac{\partial}{\partial x}\eta_{j}+\frac{\partial}{\partial y}\eta_{j})\right)\right\|_{L^{\infty}(\mathbb{R}^{2})}\leq\frac{C_{p}}{\varepsilon}\left\|\sum_{j=1}^{N}\eta_{j}\right\|_{W^{1,1}(\mathbb{R}^{2})},$
with $\widetilde{\omega}:=T\omega=(1+2\pi i(y-x))$ and by employing the Hausdorff-Young inequality in the last step. Finally, (10) can be obtained as
$\left\|A(0,\cdot)-A^{\eta}(0,\cdot)\right\|_{L^{1}(\mathbb{R})}$ $\leq\left\|\frac{\sum_{j=1}^{N}T\mathcal{F}(F_{j}-F_{j}^{\eta})}{\sum_{j=1}^{N}\mathcal{A}\phi_{j}}(0,\cdot)\right\|_{L^{1}(\Omega\cap(\{0\}\times\mathbb{R}))},$
$\leq\frac{1}{\varepsilon}\left\|\sum_{j=1}^{N}\mathcal{F}\eta_{j}(\cdot,0)\right\|_{L^{1}(\Omega_{x})},$
$\leq\frac{C_{\mathbb{R},\Omega}}{\varepsilon}\left\|\sum_{j=1}^{N}\mathcal{F}\eta_{j}(\cdot,0)\right\|_{L^{\infty}(\mathbb{R})},$
$\leq\frac{C_{\mathbb{R},\Omega}}{\varepsilon}\left\|\mathcal{F}\Big{(}\sum_{j=1}^{N}\eta_{j}\Big{)}\right\|_{L^{\infty}(\mathbb{R}^{2})},$
$\leq\frac{C_{\mathbb{R},\Omega}}{\varepsilon}\cdot\left\|\sum_{j=1}^{N}\eta_{j}\right\|_{L^{1}(\mathbb{R}^{2})},$
where the last inequality is an application of the Hausdorff-Young inequality. To bound the propagation of data noise in the reconstructions $\widetilde{f}$ and $\widetilde{f\,}^{\eta}$, we first define
$\widetilde{u}_{c}(x)$ $:=\int_{\mathbb{R}}A(x,y)e^{2\pi i(x/2+c)y}\mathrm{d}y,$
$\widetilde{u}^{\eta}_{c}(x)$ $:=\int_{\mathbb{R}}A^{\eta}(x,y)e^{2\pi i(x/2+c)y}\mathrm{d}y,$
for $c\in\mathbb{R}$ and note that their difference can be straightforwardly bounded by
$\left\|\widetilde{u}_{c}-\widetilde{u}_{c}^{\eta}\right\|_{L^{p}(\mathbb{R})}$ $\leq\left(\int_{\mathbb{R}}\left|\int_{\mathbb{R}}\left(A(x,y)-A^{\eta}(x,y)\right)e^{2\pi i(x/2+c)y}\mathrm{d}y\right|^{p}\mathrm{d}x\right)^{1/p},$
$\leq\left(\int_{\mathbb{R}}\left|\int_{\mathbb{R}}\left|A(x,y)-A^{\eta}(x,y)\right|\mathrm{d}y\right|^{p}\mathrm{d}x\right)^{1/p},$
(11) $\leq\int_{\mathbb{R}}\left(\int_{\mathbb{R}}\left|A(x,y)-A^{\eta}(x,y)\right|^{p}\mathrm{d}x\right)^{1/p}\mathrm{d}y=\left\|A-A^{\eta}\right\|_{L^{p,1}(\mathbb{R}^{2})},$
where the last inequality is an application of Minkowski’s integral inequality.
For reconstructing from $\widetilde{u}_{c},\widetilde{u}_{c}^{\eta}$, it is necessary to fix a point $c\in\mathbb{R}$ such that $|\widetilde{u}_{c}(0)|\neq 0$ and $|\widetilde{u}_{c}^{\eta}(0)|\neq 0.$ Without loss of generality and for simpler notation, suppose that one can choose $c=0$ and denote $\widetilde{u}(x):=\widetilde{u}_{0}(x),$ $\widetilde{u}^{\eta}(x):=\widetilde{u\,}^{\eta}_{0}(x).$ Then, the reconstructions are obtained by taking
$\widetilde{f}(x)$ $=\frac{1}{\sqrt{\left|\widetilde{u}(0)\right|}}\widetilde{u}(x),$
$\widetilde{f\,}^{\eta}(x)$ $=\frac{1}{\sqrt{\left|\widetilde{u}^{\eta}(0)\right|}}\widetilde{u}^{\eta}(x),$
which corresponds to setting the phase factor at $x=0$ to be equal to $1$, i.e., $\widetilde{f}(0)=|\widetilde{f}(0)|$ and $\widetilde{f\,}^{\eta}(0)=|\widetilde{f\,}^{\eta}(0)|$ Note that with the choice above, $|\widetilde{f}(0)|=\sqrt{|\widetilde{u}(0)|}$ and $|\widetilde{f\,}^{\eta}(0)|=\sqrt{|\widetilde{u}^{\eta}(0)|}$.
The propagation of data noise on the reconstructions $\widetilde{f}$ and $\widetilde{f\,}^{\eta}$ can be estimated as follows:
Proposition 3.3.
Let $k:=\sqrt{\left|\widetilde{u}(0)\right|}$ and $k^{\eta}:=\sqrt{\left|\widetilde{u}^{\eta}(0)\right|}$. Then,
(12) $\left\|\widetilde{f}-\widetilde{f\,}^{\eta}\right\|_{L^{p}(\mathbb{R})}$ $\leq\left(\frac{\left\|A\right\|_{L^{p,1}(\mathbb{R}^{2})}}{2(k\,k^{\eta})^{3/2}}+\frac{1}{k^{\eta}}\right)\frac{\max(C_{\mathbb{R},\Omega},C_{p})}{\varepsilon}\left\|\sum_{j=1}^{N}\eta_{j}\right\|_{W^{1,1}(\mathbb{R}^{2})}.$
Proof 3.4.
We have
$\left\|\widetilde{f}-\widetilde{f\,}^{\eta}\right\|_{L^{p}(\mathbb{R})}$ $\leq\left\|\frac{1}{k}\cdot\widetilde{u}-\frac{1}{k^{\eta}}\cdot\widetilde{u}\right\|_{L^{p}(\mathbb{R})}+\left\|\frac{1}{k^{\eta}}\cdot\widetilde{u}-\frac{1}{k^{\eta}}\cdot\widetilde{u}^{\eta}\right\|_{L^{p}(\mathbb{R})},$
$\leq\left|\frac{1}{k}-\frac{1}{k^{\eta}}\right|\cdot\left\|\widetilde{u}\right\|_{L^{p}(\mathbb{R})}+\frac{1}{k^{\eta}}\left\|\widetilde{u}-\widetilde{u}^{\eta}\right\|_{L^{p}(\mathbb{R})},$
$\leq\frac{1}{2(k\,k^{\eta})^{3/2}}\left|(k^{\eta})^{2}-k^{2}\right|\cdot\left\|\widetilde{u}\right\|_{L^{p}(\mathbb{R})}+\frac{1}{k^{\eta}}\left\|\widetilde{u}-\widetilde{u}^{\eta}\right\|_{L^{p}(\mathbb{R})},$
$\leq\frac{1}{2(k\,k^{\eta})^{3/2}}\left|\widetilde{u}^{\eta}(0)-\widetilde{u}(0)\right|\cdot\left\|\widetilde{u}\right\|_{L^{p}(\mathbb{R})}+\frac{1}{k^{\eta}}\left\|\widetilde{u}-\widetilde{u}^{\eta}\right\|_{L^{p}(\mathbb{R})},$
where the arithmetic-geometric mean inequality is applied in the third line (since $k,k^{\eta}>0$) and the reverse triangle inequality is used in the fourth line. Combining this with (11), we obtain
$\left\|\widetilde{f}-\widetilde{f\,}^{\eta}\right\|_{L^{p}(\mathbb{R})}\leq\frac{1}{2(k\,k^{\eta})^{3/2}}\left\|A(0,\cdot)-A^{\eta}(0,\cdot)\right\|_{L^{1}(\mathbb{R})}\cdot\left\|A\right\|_{L^{p,1}(\mathbb{R}^{2})}+\frac{1}{k^{\eta}}\left\|A-A^{\eta}\right\|_{L^{p,1}(\mathbb{R}^{2})}.$
Further estimating the RHS employing (9) and (10) yields the desired bound.
3.1.2 Approximation error
The error bound in (12) becomes small as we increase $\varepsilon$. However, this comes at the cost of increasing $\|f-\widetilde{f}\|$ since we only recover $\mathcal{A}f$ or its approximation on the bounded domain $\Omega=\bigcup_{j=1}^{N}\Omega_{j}$ decided by $\varepsilon$. The bound on the approximation error $\|f-\widetilde{f}\|$ will depend on the quantities
(13) $\left\|\mathcal{A}f\right\|_{L^{p,1}(\mathbb{R}^{2}\backslash\Omega)}:=\left\|\mathds{1}_{\mathbb{R}^{2}\backslash\Omega}\mathcal{A}f\right\|_{L^{p,1}(\mathbb{R}^{2})},$
and
(14) $\left\|\mathcal{A}f(0,\cdot)\right\|_{L^{1}((\mathbb{R}^{2}\backslash\Omega)\cap(\left\{0\right\}\times\mathbb{R}))}.$
Note that $|\mathcal{A}f|=|V_{f}f|,$ so (13) quantifies the decay of both $f$ and its Fourier transform $\widehat{f}.$ Moreover, (14) gives the decay of $|\mathcal{F}(|f|^{2})|$ since $|\mathcal{A}f(0,\cdot)|=|\mathcal{F}(|f|^{2})|$.
We further note that since $A$ and $\mathcal{A}f$ agree on $\Omega$ and $A$ is identical to zero on $\mathbb{R}^{2}\backslash\Omega$, we have
$\left\|A-\mathcal{A}f\right\|_{L^{p,1}(\mathbb{R}^{2})}=\left\|A-\mathcal{A}f\right\|_{L^{p,1}(\mathbb{R}^{2}\backslash\Omega)}=\left\|\mathcal{A}f\right\|_{L^{p,1}(\mathbb{R}^{2}\backslash\Omega)},$
and, similar to (11),
$\left\|u-\widetilde{u}\right\|_{L^{p}(\mathbb{R})}\leq\left\|\mathcal{A}f-A\right\|_{L^{p,1}(\mathbb{R}^{2})},$
where $u(x):=\int_{\mathbb{R}}\mathcal{A}f(x,y)e^{\pi ixy}\mathrm{d}y$. Note that $u(0)=|f(0)|^{2}>0$ and define $k^{\mathrm{true}}:=\sqrt{u(0)}$. As before, we take $k=\sqrt{|\widetilde{u}(0)|}$ and set the phase factors of $f$ and $\widetilde{f}$ to $1$ at $x=0$. With this, we obtain the following lemma.
Lemma 3.5.
Let $f$ be the true solution and let $\widetilde{f}$ be the reconstruction from $F_{1},\dots,F_{N}$ by formula (6). Then,
(15) $\left\|f-\widetilde{f}\right\|_{L^{p}(\mathbb{R})}\leq\frac{1}{k^{\mathrm{true}}}\left\|\mathcal{A}f\right\|_{L^{p,1}(\mathbb{R}^{2}\backslash\Omega)}+\frac{\left\|A\right\|_{L^{p,1}(\mathbb{R}^{2})}}{2(k^{\mathrm{true}}\,k)^{3/2}}\left\|\mathcal{A}f(0,\cdot)\right\|_{L^{1}((\mathbb{R}^{2}\backslash\Omega)\cap(\left\{0\right\}\times\mathbb{R}))}.$
Proof 3.6.
We first use the triangle inequality and split the approximation error into two terms. Recalling that
$\widetilde{u}(x)=\int_{\mathbb{R}}A(x,y)e^{\pi ixy}\mathrm{d}y\quad\text{and}\quad\widetilde{f}(x)=\widetilde{u}(x)/k\,,$
we obtain
$\left\|f-\widetilde{f}\right\|_{L^{p}(\mathbb{R})}$ $\leq\left\|f-\frac{\widetilde{u}}{k^{\mathrm{true}}}\right\|_{L^{p}(\mathbb{R})}+\left\|\frac{\widetilde{u}}{k^{\mathrm{true}}}-\widetilde{f}\right\|_{L^{p}(\mathbb{R})}$
$=\left\|\frac{u}{k^{\mathrm{true}}}-\frac{\widetilde{u}}{k^{\mathrm{true}}}\right\|_{L^{p}(\mathbb{R})}+\left\|\frac{\widetilde{u}}{k^{\mathrm{true}}}-\widetilde{f}\right\|_{L^{p}(\mathbb{R})}$
$\leq\frac{1}{k^{\mathrm{true}}}\left\|\mathcal{A}f\right\|_{L^{p,1}(\mathbb{R}^{2}\backslash\Omega)}+\left|\frac{1}{k^{\mathrm{true}}}-\frac{1}{k}\right|\cdot\left\|\widetilde{u}\right\|_{L^{p}(\mathbb{R})}$
$\leq\frac{1}{k^{\mathrm{true}}}\left\|\mathcal{A}f\right\|_{L^{p,1}(\mathbb{R}^{2}\backslash\Omega)}+\left|\frac{1}{k^{\mathrm{true}}}-\frac{1}{k}\right|\cdot\left\|A\right\|_{L^{p,1}(\mathbb{R}^{2})}.$
Similar to the derivations of the noise propagation term, we can further bound this by
$\left\|f-\widetilde{f}\right\|_{L^{p}(\mathbb{R})}$ $\leq\frac{1}{k^{\mathrm{true}}}\left\|\mathcal{A}f\right\|_{L^{p,1}(\mathbb{R}^{2}\backslash\Omega)}+\frac{1}{2(k^{\mathrm{true}}\,k)^{3/2}}\left|u(0)-\widetilde{u}(0)\right|\cdot\left\|A\right\|_{L^{p,1}(\mathbb{R}^{2})}$
$\leq\frac{1}{k^{\mathrm{true}}}\left\|\mathcal{A}f\right\|_{L^{p,1}(\mathbb{R}^{2}\backslash\Omega)}+\frac{1}{2(k^{\mathrm{true}}\,k)^{3/2}}\left\|\mathcal{A}f(0,\cdot)-A(0,\cdot)\right\|_{L^{1}(\mathbb{R})}\cdot\left\|A\right\|_{L^{p,1}(\mathbb{R}^{2})}$
$\leq\frac{1}{k^{\mathrm{true}}}\left\|\mathcal{A}f\right\|_{L^{p,1}(\mathbb{R}^{2}\backslash\Omega)}+\frac{1}{2(k^{\mathrm{true}}\,k)^{3/2}}\left\|\mathcal{A}f(0,\cdot)\right\|_{L^{1}((\mathbb{R}^{2}\backslash\Omega)\cap(\left\{0\right\}\times\mathbb{R}))}\cdot\left\|A\right\|_{L^{p,1}(\mathbb{R}^{2})}.$
3.1.3 Total error
We are now in a position to combine the estimates (12) and (15) on propagation of data noise and approximation error to a total error estimate.
Proposition 3.7.
For some constant $C>0$ depending on $\|A\|_{L^{p,1}(\mathbb{R}^{2})},k,k^{\eta},k^{\mathrm{true}},C_{\mathbb{R},\Omega}$ and $C_{p},$ we have
$\left\|f-\widetilde{f\,}^{\eta}\right\|_{L^{p}(\mathbb{R})}$ $\leq\frac{1}{k^{\mathrm{true}}}\left\|\mathcal{A}f\right\|_{L^{p,1}(\mathbb{R}^{2}\backslash\Omega)}+\frac{1}{2(k^{\mathrm{true}}\,k)^{3/2}}\left\|\mathcal{A}f(0,\cdot)\right\|_{L^{1}((\mathbb{R}^{2}\backslash\Omega)\cap(\left\{0\right\}\times\mathbb{R}))}\cdot\left\|A\right\|_{L^{p,1}(\mathbb{R}^{2})}$
$\quad+\left(\frac{1}{2(k\,k^{\eta})^{3/2}}\cdot\left\|A\right\|_{L^{p,1}(\mathbb{R}^{2})}+\frac{1}{k^{\eta}}\right)\cdot\frac{1}{\varepsilon}\max(C_{\mathbb{R},\Omega},C_{p})\left\|\sum_{j=1}^{N}\eta_{j}\right\|_{W^{1,1}(\mathbb{R}^{2})}$
(16) $\leq C\cdot\left(\left\|\mathcal{A}f\right\|_{L^{p,1}(\mathbb{R}^{2}\backslash\Omega)}+\left\|\mathcal{A}f(0,\cdot)\right\|_{L^{1}((\mathbb{R}^{2}\backslash\Omega)\cap(\left\{0\right\}\times\mathbb{R}))}+\frac{1}{\varepsilon}\left\|\sum_{j=1}^{N}\eta_{j}\right\|_{W^{1,1}(\mathbb{R}^{2})}\right).$
Remark 3.8.
There are three terms on the RHS of (16). The first two terms depend on $\Omega=\cup_{j=1}^{N}\Omega_{j}$, which is determined by the choice of $\varepsilon$. As $\varepsilon$ increases, so do the first two terms. Conversely, the last term is inversely proportional to $\varepsilon$. Therefore, one must balance these terms when selecting an appropriate $\varepsilon$.
Now, consider a fixed $\varepsilon$. If we only use a single window function, i.e., $N=1$, then the set $\Omega$ can be significantly smaller compared to the case with multiple windows; see Figures 1 and 2 for illustrations. This again sheds lights on the fact that we achieve better stability by using more windows. In this way, we can significantly reduce the magnitude of the first two terms in (16) while keeping the last term constant.
3.2 Stability of Algorithm 2
In Algorithm 2, the reconstruction is done separately on each $\Omega_{j}$, using $F_{j}$, $F_{j}^{\eta}$ for computing $A$, $A^{\eta}$, respectively. Again, $A$ and $A^{\eta}$ will be set to zero outside of $\Omega$ and we obtain the following error bound for the data noise propagation:
Lemma 3.9.
Let $A,A^{\eta}$ be the reconstructions obtained from $F_{j},$ $F_{j}^{\eta},$ respectively, $j=1,\dots,N$, and for some $\varepsilon>0$, let $\Omega_{j}$ be chosen such that $\left|\mathcal{A}h_{n_{j}}(x,y)\right|>\varepsilon$ for all $(x,y)\in\Omega_{j}$ for all $j\in\{1,\dots,N\}$. Then,
$\left\|A-A^{\eta}\right\|_{L^{p,1}(\mathbb{R}^{2})}$$\leq$ $\frac{C_{p}}{\varepsilon}\max_{j=1,\dots,N}\left\|\eta_{j}\right\|_{W^{1,1}(\mathbb{R}^{2})},$
$\left\|A(0,\cdot)-A^{\eta}(0,\cdot)\right\|_{L^{1}(\mathbb{R})}$$\leq$ $\frac{C_{\mathbb{R},\Omega}}{\varepsilon}\cdot\max_{j=1,\dots,N}\left\|\eta_{j}\right\|_{L^{1}(\mathbb{R}^{2})}.$
Proof 3.10.
By the reconstruction formula,
$\left\|A-A^{\eta}\right\|_{L^{p,1}(\mathbb{R}^{2})}\leq\sum_{j=1}^{N}\left\|\mathds{1}_{\Omega_{j}}\frac{T\mathcal{F}(F_{j}-F_{j}^{\eta})}{\mathcal{A}h_{n_{j}}}\right\|_{L^{p,1}(\mathbb{R}^{2})}\leq\sum_{j=1}^{N}C_{j,p}\cdot\left\|\mathds{1}_{\Omega_{j}}\omega\frac{T\mathcal{F}(F_{j}-F_{j}^{\eta})}{\mathcal{A}h_{n_{j}}}\right\|_{L^{\infty}(\mathbb{R}^{2})},$
where $C_{j,p}:=\|\mathds{1}_{\Omega_{j}}/\omega\|_{L^{p,1}(\mathbb{R}^{2})}<\infty$ and instead of (9), we now have the similar bound
$\left\|A-A^{\eta}\right\|_{L^{p,1}(\mathbb{R}^{2})}$ $\leq\sum_{j=1}^{N}\frac{C_{j,p}}{\inf_{\Omega_{j}}|\mathcal{A}h_{n_{j}}|}\cdot\left\|\widetilde{\omega}\,\mathcal{F}\eta_{j}\right\|_{L^{\infty}(\mathbb{R}^{2})}$
$\leq\frac{1}{\varepsilon}\sum_{j=1}^{N}C_{j,p}\cdot\left\|\mathcal{F}\left(\eta_{j}-\frac{\partial}{\partial x}\eta_{j}+\frac{\partial}{\partial y}\eta_{j}\right)\right\|_{L^{\infty}(\mathbb{R}^{2})}$
$\leq\frac{1}{\varepsilon}\sum_{j=1}^{N}C_{j,p}\cdot\left\|\eta_{j}\right\|_{W^{1,1}(\mathbb{R}^{2})}\leq\frac{C_{p}}{\varepsilon}\max_{j=1,\dots,N}\left\|\eta_{j}\right\|_{W^{1,1}(\mathbb{R}^{2})},$
where we have used $\sum_{j=1}^{N}C_{j,p}=C_{p}$ since $\Omega_{1},\dots,\Omega_{N}$ are pairwise disjoint and the Hausdorff-Young inequality. Similar to (10), we can further derive
$\left\|A(0,\cdot)-A^{\eta}(0,\cdot)\right\|_{L^{1}(\mathbb{R})}$ $\leq\sum_{j=1}^{N}\left\|\frac{T\mathcal{F}(F_{j}-F_{j}^{\eta})}{\mathcal{A}h_{n_{j}}}(0,\cdot)\right\|_{L^{1}(\Omega_{j}\cap(\{0\}\times\mathbb{R}))}$
$\leq\frac{1}{\varepsilon}\sum_{j=1}^{N}\left\|\mathcal{F}\eta_{j}(\cdot,0)\right\|_{L^{1}(\Omega_{j,x})}$
$\leq\frac{1}{\varepsilon}\sum_{j=1}^{N}\left\|\mathcal{F}\eta_{j}(\cdot,0)\right\|_{L^{\infty}(\mathbb{R})}\cdot\left|\Omega_{j,x}\right|$
$\leq\frac{1}{\varepsilon}\sum_{j=1}^{N}\left\|\eta_{j}\right\|_{L^{1}(\mathbb{R}^{2})}\cdot\left|\Omega_{j,x}\right|$
$\leq\frac{C_{\mathbb{R},\Omega}}{\varepsilon}\cdot\max_{j=1,\dots,N}\left\|\eta_{j}\right\|_{L^{1}(\mathbb{R}^{2})},$
where $\Omega_{j,x}=\{x\in\mathbb{R}:(x,0)\in\Omega_{j}\}\subset\mathbb{R}$.
With this, the propagation of data noise can be estimated by
$\left\|\widetilde{f}-\widetilde{f\,}^{\eta}\right\|_{L^{p}(\mathbb{R})}$ $\leq\left(\frac{1}{2(k\,k^{\eta})^{3/2}}\cdot\left\|A\right\|_{L^{p,1}(\mathbb{R}^{2})}+\frac{1}{k^{\eta}}\right)\cdot\frac{1}{\varepsilon}\max(C_{\mathbb{R},\Omega},C_{p})\max_{j=1,\dots,N}\left\|\eta_{j}\right\|_{W^{1,1}(\mathbb{R}^{2})}.$
The approximation error is the same as in the previous case (cf. (15)), so that the total error can be bounded by
$\left\|f-\widetilde{f\,}^{\eta}\right\|_{L^{p}(\mathbb{R})}$ $\leq\frac{1}{k^{\mathrm{true}}}\left\|\mathcal{A}f\right\|_{L^{p,1}(\mathbb{R}^{2}\backslash\Omega)}+\frac{1}{2(k^{\mathrm{true}}\,k)^{3/2}}\left\|\mathcal{A}f(0,\cdot)\right\|_{L^{1}((\mathbb{R}^{2}\backslash\Omega)\cap(\left\{0\right\}\times\mathbb{R}))}\cdot\left\|A\right\|_{L^{p,1}(\mathbb{R}^{2})}$
$\quad+\left(\frac{1}{2(k\,k^{\eta})^{3/2}}\cdot\left\|A\right\|_{L^{p,1}(\mathbb{R}^{2})}+\frac{1}{k^{\eta}}\right)\cdot\frac{1}{\varepsilon}\max(C_{\mathbb{R},\Omega},C_{p})\max_{j=1,\dots,N}\left\|\eta_{j}\right\|_{W^{1,1}(\mathbb{R}^{2})}$
(17) $\leq C\cdot\left(\left\|\mathcal{A}f\right\|_{L^{p,1}(\mathbb{R}^{2}\backslash\Omega)}+\left\|\mathcal{A}f(0,\cdot)\right\|_{L^{1}((\mathbb{R}^{2}\backslash\Omega)\cap(\left\{0\right\}\times\mathbb{R}))}+\frac{1}{\varepsilon}\max_{j=1,\dots,N}\left\|\eta_{j}\right\|_{W^{1,1}(\mathbb{R}^{2})}\right),$
for some constant $C>0$ depending on $\|A\|_{L^{p,1}(\mathbb{R}^{2})},k,k^{\eta},k^{\mathrm{true}},C_{\mathbb{R},\Omega}$ and $C_{p}.$
The third term on the RHS is the only difference between (16) and (17). The former depends on the maximum noise power, while the latter involves the sum of the noise for different window functions. Their differences result from the distinctions between Algorithms 1 and 2 regarding how the multi-window measurements are utilized in the STFT phase retrieval.
4 Numerical examples
In this section, we present a few 1D phase retrieval examples using the explicit formula (2)-(4), given the phaseless measurement $|V_{g}f|$ and the ambiguity function $\mathcal{A}g$ of the window function $g$ or $V_{g}g$.
4.1 Discretization
Recall the continuous STFT transform given in Equation (1). If we know the function $f$ is compactly supported on a finite interval of $\mathbb{R}$, e.g., $[-T,T]$ for some $T>0$, then we can perform the integration on $[-T,T]$:
(18) $V_{g}f(x,y):=\int_{-T}^{T}f(t)\overline{g(t-x)}e^{-2\pi ity}\,\mathrm{d}t.$
Without loss of generality, we only consider periodic window functions $g(t)$ on $[-T,T]$, i.e., $g(-t)=g(2T-t)$. All Hermite and Gauss windows used in the numerical examples are periodized.
For the numerical integration, consider an equidistant discretization in time such that $f_{l}:=f(x_{l})$, $x_{l}=-T+l\Delta t$, $10\leq l\leq L-$ 1 and $L=2T/\Delta t$. We compute the STFT at $K$ equidistant time-domain shifts $x_{0},\ldots,x_{k},\ldots,x_{K-1}$, based on the uniform spacing $a\Delta t$ with $a\in\mathbb{N}^{+}$ and $K=L/a$. Similarly, we discretize the frequency domain interval $\left[-\frac{1}{2\Delta t},\frac{1}{2\Delta t}\right]$ uniformly by taking $M$ points, i.e., $y_{m}=\frac{1}{\Delta t}\left(-\frac{1}{2}+\frac{m}{M}\right)$ for $0\leq m<M$. Note that $1/(2\Delta t)$ is the Nyquist frequency. Using the left Riemann sum, (18) can be approximated by
(19) $c_{\left(m,k\right)}=\Delta t\sum_{l=0}^{L-1}f_{l}\,\overline{g_{l-ka+L/2}}\,e^{-2\pi i(-T+l\Delta t)\frac{1}{\Delta t}\left(-\frac{1}{2}+\frac{m}{M}\right)},$
where $m=0,\ldots,M-1,$ and $k=0,\ldots,K-1,$ with
$c_{\left(m,k\right)}\approx V_{g}f\left(-T+ka\Delta t,\frac{1}{\Delta t}\left(-\frac{1}{2}+\frac{m}{M}\right)\right)$
being the $(m+1,k+1)$-th entry of the coefficient matrix $c$. The index $l-ka+L/2$ in (19) is computed modulo $L$ since we assume that the window function $g$ is periodic. That is,
$g_{-s}=g(-T-s\Delta t)=g(T-s\Delta t)=g_{(2T-s\Delta t)/\Delta t}=g_{L-s},\quad\forall\,s=0,\ldots,L-1.$
There are many fast ways of implementing (19); see the toolbox [23] for an example. One should note that if $\Delta t=1$, (19) reduces to the so-called Discrete Short-Time Fourier Transform (Discrete STFT).
Since the output of our proposed algorithms recovers the original signal up to a global phase, we define the following misfit function $d(\,\cdot\,,\,\cdot\,)$ that measures the relative error between two signals $\text{f}_{1}$ and $\text{f}_{2}$ up to a global phase factor:
(20) $d(\text{f}_{1},\text{f}_{2})=\min_{\theta\in[0,2\pi)}\frac{\|\text{f}_{1}\ -e^{i\theta}\text{f}_{2}\|_{\ell^{2}}}{\|\text{f}_{1}\|_{\ell^{2}}}.$
The angle $\theta^{*}$ that minimizes (20) is also the missing global phase angle.
(a) Chirp signal $f(t)$
(b) $\log|\mathcal{A}f(x,y)|$
Figure 3: (a) The ground truth chirp signal $f(t)$; (b) The absolute value of the ambiguity function of $f$ plotted in the log scale.
(a) Recovery with $40$ FRFT dilated Gauss windows
(b) Recovery with a standard Gauss window
Figure 4: (a) Recovery using $40$ fractional Fourier transform of the dilated Gauss windows at various angles; (b) Recovery using a single standard Gauss window. The ground truth is plotted in blue.
(a) Using $6$ Hermite windows picked uniformly
(b) Using $6$ Hermite windows picked randomly
Figure 5: (a) Recovery using $6$ Hermite windows of $h_{n}$ where $n=0,10,20,30,40,50$; (b) Recovery using $6$ Hermite windows of $h_{n}$ with the index $n$ randomly selected between 00 and $50$, resulting in a set of indices $\{17,24,29,30,31,49\}$. The ground truth is plotted in blue.
4.2 Synthetic examples
In this subsection, we will present two STFT phase retrieval examples using the proposed multi-window Algorithms 1 and 2.
4.2.1 Chirp signal recovery
As our first inversion example, we consider the chirp signal shown in Figure 3a. As illustrated in Figure 3b, its ambiguity function has extensive support in the time-frequency domain, making it challenging to reconstruct using a single window of the standard Gauss function whose ambiguity function has a very small essential support as seen in Figure 1a. We set $\varepsilon=10^{-3}$ as a cutoff value, as used in Algorithms 1 and 2 to prevent the instability from dividing the phaseless measurements by a small value.
(a) Gaussian mixture signal $f(t)$
(b) $\log|\mathcal{A}f(x,y)|$
Figure 6: Multi-modal function example: (a) The ground truth $f(t)$; (b) The absolute value of the ambiguity function of $f$ plotted in the log scale.
We first use $40$ fractionally transformed dilated Gaussian windows to perform the reconstruction following Algorithm 1, i.e., $\mathcal{F}_{\alpha_{j}}\varphi^{a}$ with $a=15$ and $\alpha_{j}=(j-1)\pi/40$, $j=1,\ldots,40$. The relative error based on formula (20) between the ground truth and the reconstructed signal is $0.0084$. We also use the standard Gauss function $\varphi^{1}$ as a single window function for comparison. The relative error is $0.7928$. After finding the global phase factor, we plot the reconstructions under different measuring window schemes in Figure 4. More measuring windows yield a much more extensive joint coverage of the time-frequency domain, resulting in better signal reconstruction.
Next, we use multiple Hermite windows to perform the recovery. Consider $6$ Hermite function $h_{n}$ with $n\in\{0,10,20,30,40,50\}$. We then follow Algorithm 2 to perform the inversion. The misfit between the reconstructed signal and the ground truth measured by (20) is $0.0025$. On the other hand, we want to test the idea of randomly choosing window functions. We draw $6$ windows uniformly from the zeroth to the fiftieth degree Hermite functions, and use them to generate the phaseless data and perform the STFT phase retrieval following Algorithm 2. We repeat this procedure $100$ times to study the performance of this inversion algorithm. The mean relative error measured by (20) between the ground truth and the reconstructed signal is $0.0298$. In particular, $90\%$ of the randomly picked window sets yield signal reconstructions with relative error less than $0.0320$. We also plot the reconstruction results from the uniformly chosen windows and one realization of the randomly selected windows in Figure 5.
4.2.2 Recovery of a multi-modal function with noisy measurements
Figure 7: Multi-modal function example: the noise-free measurement of $|V_{g}f|$ for $g=\varphi^{1}$ (left) and the noisy data used in phase retrieval (right).
We consider a multi-modal function plotted in Figure 6a, a linear combination of time-shifts of $\varphi^{4}$, the dilated Gauss function. The absolute value of its ambiguity function in log scale is shown in Figure 6b. This signal is known to cause stability issues in STFT phase retrieval [2, Sec. 6]. The measurement is contaminated by a multiplicative noise following the normal distribution $\mathcal{N}(1,0.05^{2})$; see Figure 7 for an example.
We consider the cutoff value $\varepsilon=0.5$ to define the essential support of the ambiguity function in Algorithms 1 and 2. Since $\varepsilon$ is large, we need to measure windows whose ambiguity functions have broader support in the time-frequency domain to perform the inversion. We first use fractionally Fourier-transformed versions of the dilated Gauss function $\varphi^{50}$ as the measuring windows. Since the essential support of these ambiguity functions is thin and long, we need to take more windows at different angles to achieve broad coverage of the joint essential support in the time-frequency domain: we use $80$ angles with $\alpha_{j}=(j-1)\pi/40$, $j=1,\ldots,80$. The recovery is shown in Figure 8, compared with the case using the standard Gauss as the measuring function. As seen in Figure 6b, the ambiguity function of the ground truth has almost disjoint support in the time-frequency domain. It thus cannot be stably recovered by a single Gauss window, as shown in Figure 8b, while the multiple fractional window strategy provides a stable recovery even under noisy measurement; see Figure 8a.
(a) Recovery with $80$ FRFT dilated Gaussian windows
(b) Recovery with a standard Gauss window
Figure 8: Multi-modal function example: (a) Recovery using $80$ fractional Fourier transforms of the dilated Gauss window at various angles; (b) Recovery using a single standard Gauss window. The ground truth is plotted in blue.
Next, we also use Hermite functions as the measuring windows. In Figure 9a, we use $h_{50}$ and $h_{100}$, the Hermite functions of degree $50$ and $100$, to perform STFT phase retrieval following Algorithm 2. We recover most features of the multi-modal function with minimal impact from the measurement noise illustrated in Figure 7. We also compare the inversion result using the single window $h_{100}$ in Figure 9b. Although $\mathcal{A}h_{100}$ has extensive overall coverage in the time-frequency domain, it also has $100$ circular rings that are excluded since their absolute value is below the cutoff value $\varepsilon$. As a result, the reconstruction is affected by those missing data. The ambiguity function $\mathcal{A}h_{50}$ of the additional window $h_{50}$ makes up most of those “missing” rings, so the reconstruction is much better.
The large cutoff value $\varepsilon$ mitigates the impact of noise, while using multiple windows improves the coverage in the time-frequency domain of the joint essential support of the corresponding ambiguity functions. This is evident from both using multiple fractionally Fourier-transformed dilated Gauss functions and Hermite functions of various degrees.
(a) Using $2$ Hermite windows $h_{50}$ and $h_{100}$
(b) Using a single Hermite window $h_{100}$
Figure 9: Multi-modal function example: (a) Recovery using two Hermite windows $h_{n}$ where $n=50,100$, and the relative error is $0.0899$; (b) Recovery using the single Hermite window $h_{100}$, and the relative error is $0.4403$ measured by (20). The true signal is plotted in blue.
4.3 Audio data example
Our final numerical example involves recovering a piece of an audio signal, as illustrated in Figure 10. This audio signal is significantly more complex than the earlier synthetic examples, yet the proposed multi-window strategy proves effective for the phase retrieval task.
We utilize Algorithm 2 with different sets of Hermite functions. The resulting audio files, saved as .wav files, are provided in the supplementary materials. To visualize the reconstruction performance, we plot a small segment of both the true and recovered signals. The true audio signal is shown in Figure 11a, while the recovered segments are displayed in Figures 11b, 11c, and 11d, corresponding to the use of the window function $h_{1}$, the pair $\{h_{1},h_{2}\}$, and the set $\{h_{0},h_{5},h_{10},h_{15},h_{20}\}$, respectively. As more windows are incorporated according to Algorithm 2, the quality of the audio signal recovery progressively improves.
Figure 10: The reference audio signal (in full length).
(a) True signal segment
(b) Recover with Hermite window $h_{1}$
(c) Recover with Hermite windows $h_{1},h_{2}$
(d) Recover with Hermite windows $\{h_{0},h_{5},h_{10},h_{15},h_{20}\}$
Figure 11: Audio data STFT phase retrieval example: (a) The true signal segment; (b) Recovery using single Hermite window $h_{1}$; (c) Recovery using two Hermite windows $h_{1}$ and $h_{2}$; (d) Recovery using five Hermite windows $\{h_{0},h_{5},h_{10},h_{15},h_{20}\}$.
5 Conclusions
In this paper, we introduce two multi-window strategies to improve the stability of STFT phase retrieval. The first approach uses fractional Fourier transforms of dilated Gaussian functions at various angles, while the second employs Hermite functions of different degrees as window functions. As proven in our theory, both methods effectively broaden the coverage of the ambiguity function in the time-frequency domain, significantly mitigating the instability issues associated with traditional single-window approaches. Our numerical experiments demonstrate that these multi-window strategies greatly improve the stability and accuracy of phase retrieval, even for complex audio signals. These results suggest that combining measurements from carefully chosen windows can lead to more robust and reliable signal reconstruction through our simple and direct Algorithms 1 and 2.